# Rethinking Domain-Specific LLM Benchmark Construction: A Comprehensiveness-Compactness Approach

### v2

@@ 5.5 Ablation Study of Comp-Comp
First, the benchmark optimization threshold td achieves unprecedented evaluation efficiency through distribution-aware question selection. When configured at td=0.1, the framework requires only 1.7% of benchmark questions ($Q$) while exposing conventional benchmarks’ inherent distributional biases—models show a moderate 13.5% performance drop compared to full benchmarks (Table 4). This performance gap reveals how traditional approaches overrepresent questions where models have localized competence, creating artificial inflation of aggregate scores. Comp-Comp counters this through adaptive threshold adjustments that enforce uniform question distribution across knowledge domains, rectifying evaluation skewness while maintaining testing integrity.
@@
Third, the synergistic operation of td and tc establishes a new efficiency frontier in evaluation system design. As empirically validated in Table 4, it attains performance parity with conventional benchmarks while using only 1.7% of $Q$ and 46.4% of $C$. This 98.3% reduction in questions and 53.6% decrease in corpus components demonstrates Comp-Comp’s ability to decouple evaluation quality from resource quantity, a critical advancement for sustainable benchmarking. The framework’s adaptive integration mechanism dynamically balances these thresholds, ensuring comprehensive domain coverage despite drastic resource reductions.
@@ 6 Conclusion
This work proposes Comp-Comp, a novel domain-specific LLM benchmarking framework using comprehensiveness and compactness as the guiding principle, which emphasizes the crucial role of corpora-question interaction. We developed XUBench, a large-scale benchmark in a closed academic domain, which demonstrates the effectiveness of Comp-Comp. This work serves as a valuable reference for future studies, highlighting the importance of precise dataset curation in enhancing domain-specific LLM performance.
@@
This paper proposed a case study on the academic domain to apply the Comp-Comp framework for benchmark construction. However, Comp-Comp is a general adaptable technique that can also be used in other domains such as medicine, law, finance, etc. We will explore this further integration in future work and engage in more comprehensive experiments for the research community.
@@ Appendix A Experiment Results
For the in-context learning experiments, we evaluate XUBench on several backbone models, including Vicuna-7B, Vicuna-13B Chiang et al. (2023), LLaMA3-8B Touvron et al. (2023), GLM4-9B GLM et al. (2024), Qwen2-7B Bai et al. (2023), GPT3.5, and GPT4 OpenAI (2024). Each of them is tested on three prompt templates: vanilla, few-shot, and RAG. For the RAG framework, we embed the external knowledge $Cor$ with Gte-Large Li et al. (2023) encoder and select top-k similar chunks as the context, with $k$ being 5.
@@
Table 5: The performance of In-Context Learning (IC) approaches (i.e., few-shot learning and RAG) on several question sets of XUBench with Frozen Models (FZ).
@@
Table 6: This table demonstrates the XUBench results in a Fine-Tuning (FT) based approach.
Appendix B XUBench Construction Details
@@ B.1 Question Format
To analyze the effects of the balance in precision and recall, XUBench incorporates various question formats:

### v3

@@ 5.5 Ablation Study of Comp-Comp
First, the benchmark optimization threshold td achieves unprecedented evaluation efficiency through distribution-aware question selection. When configured at td=0.1, the framework requires only 1.7% of benchmark questions ($Q$) while exposing conventional benchmarks’ inherent distributional biases, models show a moderate 13.5% performance drop compared to full benchmarks (Table 4). This performance gap reveals how traditional approaches overrepresent questions where models have localized competence, creating artificial inflation of aggregate scores. Comp-Comp counters this through adaptive threshold adjustments that enforce uniform question distribution across knowledge domains, rectifying evaluation skewness while maintaining testing integrity.
@@
Third, the synergistic operation of td and tc establishes a new efficiency frontier in the evaluation system. As validated in Table 4, it attains performance parity with conventional benchmarks while using only 1.7% of $Q$ and 46.4% of $C$. This 98.3% reduction in questions and 53.6% decrease in corpus demonstrate Comp-Comp’s ability to decouple evaluation quality from resource quantity, a critical advancement for sustainable benchmarking. Specifically, 80% of experimental trials demonstrated superior scores compared to using the original corpus. This paradigm shift suggests that conventional corpora contain substantial knowledge redundancy, overlapping or marginally informative entries that dilute retrieval effectiveness.
5.6 User Survey
We conducted a user survey with five domain experts to evaluate 100 randomly sampled LLM responses on Helpfulness, Conciseness, and Correctness using a 1-5 Likert scale, as shown in Table 5.
A key finding is the discrepancy between human evaluations and automated metrics. For frozen models (LLaMA3-8B, GLM4-9B), human evaluators preferred responses from vanilla and few-shot methods. Conversely, the RAG method, despite lower human scores, consistently achieved the highest BLEU scores. This trend persists for fine-tuned (FT) models, where RAG again leads in BLEU scores while not always being preferred by humans. Notably, all human scores for FT models were lower than their frozen counterparts. This highlights a limitation of relying solely on automated metrics like BLEU, which measure textual overlap but may not capture user-perceived quality such as coherence or relevance. Therefore, human evaluation is crucial for validating benchmark quality and guiding the development of genuinely user-centric models. These insights reinforce the necessity of a multi-faceted evaluation approach, where automated scores are always contextualized with human-centric assessments to build models that are not only accurate but also genuinely useful.
5.7 Application of Comp-Comp On More Domains
To test its generalizability, we applied the Comp-Comp framework to four benchmarks in the medical (MedQA Jin et al. (2021), PubMedQA Jin et al. (2019)) and finance (FiQA Shah et al. (2022), FinanceBench Islam et al. (2023)) domains. We analyzed the impact of various thresholds $t_{d}$ (for questions $Q$) and $t_{c}$ (for corpora $C$) to observe their performance metrics on LLaMA3.1-8B-Instruct, respectively. The results are shown in Table 6.
The results consistently show that filtering creates smaller, more efficient datasets without degrading, and often improving, performance. For instance, on MedQA, applying a threshold of $t_{d}=0.5$ reduced the number of questions from 1273 to just 28, yet increased accuracy from 0.684 to 0.786. Similarly, filtering the corpus with $t_{c}$ maintained or enhanced accuracy while using a more compact set of documents. This trend holds in the finance domain, where metrics remained stable or improved despite data reduction. Even on the difficult FinanceBench dataset, where baseline performance was low, our framework did not hurt performance.
These findings validate that Comp-Comp is an efficient and broadly applicable framework. It successfully builds compact yet comprehensive benchmarks by removing redundancy while preserving evaluation integrity, demonstrating its value for creating robust, domain-specific evaluation tools.
@@ 6 Conclusion
This work proposes Comp-Comp, which is a novel domain-specific LLM benchmarking framework using comprehensiveness and compactness as the guiding principles. To validate its performance, we developed PolyBench, a large-scale benchmark in a closed academic domain, demonstrating the effectiveness of Comp-Comp. This work serves as a valuable reference for future studies, highlighting the importance of precise dataset curation in enhancing domain-specific LLMs.
@@
Although Comp-Comp successfully demonstrates its utility within an academic setting, where the domain knowledge is well-structured and relatively static. However, more dynamic domains may pose unique challenges in terms of data availability and rapidly evolving knowledge structures that could influence the framework’s performance. Future work should extend our experiments to a broader set of domains to rigorously assess the generalizability and robustness of the COMP-COMP principles.
Furthermore, the current implementation of COMP-COMP is best suited for "closed" or well-defined domains where a comprehensive body of knowledge can be collected and stabilized. The framework’s iterative expansion process assumes a target knowledge space $S$ that is largely finite. Its applicability to "open" and continuously evolving domains, such as news analysis or social media trends, is less clear. Adapting the COMP-COMP framework to handle such dynamic data presents a significant avenue for future research.
@@
Acknowledgements
The research described in this paper has been partly supported by National Natural Science Foundation of China (Grant No.: 62372314), Hong Kong Research Grants Council under the General Research Fund (project no. PolyU15200023), a Collaborative Research Fund (project no. C1043-24G), as well as internal research funds from The Hong Kong Polytechnic University (project no. P0052406 and P0052986). This work was also supported by computational resources provided by The Centre for Large AI Models (CLAIM) of The Hong Kong Polytechnic University.
@@ Appendix A Experiment Results
For the in-context learning experiments, we evaluate PolyBench on several backbone models, including Vicuna-7B, Vicuna-13B Chiang et al. (2023), LLaMA3-8B Touvron et al. (2023), GLM4-9B GLM et al. (2024), Qwen2-7B Bai et al. (2023), GPT3.5, and GPT4 OpenAI (2024). Each of them is tested on three prompt templates: vanilla, few-shot, and RAG. For the RAG framework, we embed the external knowledge $Cor$ with Gte-Large Li et al. (2023) encoder and select top-k similar chunks as the context, with $k$ being 5.
@@
Table 7: The performance of In-Context Learning (IC) approaches (i.e., few-shot learning and RAG) on several question sets of PolyBench with Frozen Models (FZ).
@@
Table 8: This table demonstrates the PolyBench results in a Fine-Tuning (FT) based approach.
Appendix B PolyBench Construction Details
@@ B.1 Question Format
To analyze the effects of the balance in precision and recall, PolyBench incorporates various question formats: